HETAV PATEL

Grade 11,

What is area of plates of 3F parallel plates capacitor,if the seperation between the plates is 3mm?

What is area of plates of 3F parallel plates capacitor,if the seperation between the plates is 3mm?

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Capacitance of a parallel plate capacitor is given as-

C=\dfrac{A\epsilon_0}{d}
\implies A=\dfrac{Cd}{\epsilon_0}=1.017\times 10^{9}m^2